# Adversarial Bug Reports as a Security Risk in Language Model-Based Automated Program Repair

### v1

@@ Abstract.
Large Language Model (LLM) - based Automated Program Repair (APR) systems are increasingly integrated into modern software development workflows, offering automated patches in response to natural language bug reports. However, this reliance on untrusted user input introduces a novel and underexplored attack surface. In this paper, we investigate the security risks posed by adversarial bug reports—realistic-looking issue submissions crafted to mislead APR systems into producing insecure or harmful code changes. We develop a comprehensive threat model and conduct an empirical study to evaluate the vulnerability of state-of-the-art APR systems to such attacks. Our demonstration comprises 51 adversarial bug reports generated across a spectrum of strategies, from manual curation to fully automated pipelines. We test these against leading APR model and assess both pre-repair defenses (e.g., LlamaGuard variants, PromptGuard variants, Granite-Guardian, and custom LLM filters) and post-repair detectors (GitHub Copilot, CodeQL). Our findings show that current defenses are insufficient: 90% of crafted bug reports triggered attacker-aligned patches. The best pre-repair filter blocked only 47%, while post-repair analysis—often requiring human oversight—was effective in just 58% of cases. To support scalable security testing, we introduce a prototype framework for automating the generation of adversarial bug reports. Our analysis exposes a structural asymmetry: generating adversarial inputs is inexpensive, while detecting or mitigating them remains costly and error-prone. We conclude with practical recommendations for improving the robustness of APR systems against adversarial misuse and highlight directions for future work on trustworthy automated repair.
automated program repair, security
$c$s: Software and its engineering Automatic programmingccs: Security and privacy Software security engineering
@@ 1. Introduction
Automated Program Repair has emerged as a promising solution to reduce developers’ effort in maintaining large software systems. By leveraging advances in machine learning and code synthesis, APR systems automatically generate candidate patches for bugs reported by users or detected by automated testing. This capability has led to widespread interest in integrating APR into continuous integration (CI) pipelines and developer workflows.
However, as APR becomes more pervasive, its security implications remain insufficiently explored. Existing research on APR primarily focuses on patch correctness, plausibility, or overfitting, but largely overlooks the possibility that malicious actors could exploit the repair process itself as an attack vector. In particular, the reliance of APR systems on external bug reports—often submitted through open channels with minimal vetting—introduces a new attack surface that has not been systematically studied.
@@
To answer these questions, we systematically design attacks that exploit common APR workflows. We conduct experiments using state-of-the-art language models and representative open-source software repositories. Our study measures the feasibility, success rates, and cost implications of such attacks, and discusses possible hardening strategies for APR systems and project workflows.
By highlighting a novel class of threats in the software supply chain, our work urges the community to reconsider long-held trust assumptions in automated repair pipelines. We advocate for a security-aware design of APR systems—one that balances repair effectiveness with resilience to adversarial misuse.
This is no longer a purely theoretical concern: recent real-world attacks have exploited issue tracking systems to inject malicious inputs into downstream automation. For instance, adversaries have used crafted Jira tickets to trigger AI-driven actions as part of a multi-stage attack on Atlassian’s MCP infrastructure (French, 2025; Waizel et al., 2025), and prompt injection vulnerabilities have been demonstrated in GitLab’s AI assistant (Willison, [n. d.]). These incidents underscore the urgent need to secure the interfaces between user-submitted bug reports and automated code generation systems.
1.1. Contributions
1. Threat Model: A formal model describing how malicious actors can exploit bug reporting interfaces to subvert APR.
2. Attack Framework: We provide prompts used for attacks as well as an open-source prototype that automates the generation and evaluation of adversarial bug reports.
3. Empirical Study of Defenses: A systematic evaluation of pre- and post-APR defenses, including LLM-based filters and static analysis tools, revealing key limitations and trade-offs.
1.2. Replication
A complete replication package, including automation scripts for generating malicious bug reports, code to orchestrate the attack workflows, all defense mechanisms with prompts, as well as the full set of issues, generated patches, and execution trajectories, is available at: https://anonymous.4open.science/r/apr-sec-anon-1486/.
@@ 2.1. Background and Threat Landscape
However, this design introduces a novel attack surface. LLM-driven APR systems inherently trust user-submitted bug reports and lack rigorous safeguards to verify intent, semantic correctness, or security implications of proposed changes. This makes them vulnerable to adversarial bug reports—crafted inputs that appear plausible but manipulate the APR pipeline into producing insecure or incorrect patches.
The security implications of LLM-based developer tools have recently gained attention, especially through the OWASP Top 10 for LLM Applications (Wilson and Dawson, 2023), which highlights prompt injection, overreliance, and insecure output handling as critical risks. Although defenses such as LlamaGuard (Inan et al., 2023), PromptGuard (Meta, 2025), and CodeQL (Youn et al., 2023) exist, they are not yet integrated into typical APR workflows. Moreover, most APR tools are not hardened against misuse of their input channels.
Despite growing awareness of security issues in LLM applications, vulnerabilities in APR workflows remain underexplored. Our work builds on this gap, investigating how adversarial bug reports can exploit LLM-powered repair systems, compromise CI/CD pipelines, and waste developer and compute resources with minimal effort.
@@ 2.2. End-to-End APR Pipeline and Attack Walkthrough
Figure 1 illustrates a standard APR pipeline commonly used in modern software projects. The process begins when a user submits a bug report through the issue tracker. In our illustrative attack scenario, the adversary submits a malicious issue designed to trigger a reversion of a previously fixed CVE (Figure 2(a)). The issue appears plausible and well-formed, making it likely to pass initial filtering.
Typically, the bug report passes through automated triage mechanisms. For APR these may include: a security filter, such as LlamaGuard4, which assesses the textual input for prompt injections or unsafe patterns; and an APR suitability filter, which checks whether the report is appropriate for repair using a LLM with a structured prompt. Here, the issue is incorrectly classified as both non-malicious (LlamaGuard4) and valid for APR (GPT-4.1-mini).
Once accepted, the APR system attempts to reproduce the defect, often within a containerized testbed. This step may involve compiling the code, executing test cases, or generating synthetic inputs. Upon successful reproduction, a candidate patch is synthesized using templates, retrieval-based methods, or generative models. In our example, the patch effectively reverts the original CVE fix (Figure 2(b)).
Next, the patch undergoes validation via automated test suites in a CI/CD environment. Since test coverage is rarely complete—especially for security fixes—such regressions often go undetected. If the patch passes, a pull request is automatically created. Mind, that CI/CD environments may be also a target for the attacker (but not in running example).
This PR may be reviewed automatically (e.g., by GitHub Copilot, Figure 2(c)) and scanned by static analysis tools like CodeQL (Figure 2(d)). However, these tools may fail to detect subtle regressions or logic flaws introduced by the adversary as occurred in the presented example.
Manual review can help, but only if reviewers are involved early, have sufficient expertise, and invest the effort to investigate. In practice, limited context and reviewer fatigue make it easy for well-crafted reports and plausible code to slip through.
@@ 2.3. Threat Model
Source Code Integrity and security must be maintained—no backdoors or reintroduced vulnerabilities.
@@
Impact APR and CI/CD environment (DOS, runtime exploits, info leaks).
Source code (vulnerable or malicious patches may bypass review and reach production.).
@@
(Spoofing and Repudiation less relevant in open bug reporting.)
To assess the security risks posed by adversarial bug reports, we define a structured threat model that captures both the system’s critical assets and the adversary’s capabilities. Our focus is on language model-driven APR pipelines, which combine natural language inputs, automated patch generation, and integration into CI/CD workflows.
@@ 3. Adversarial Attack Strategies (RQ1)
To assess the susceptibility of APR systems to adversarial manipulation (RQ1), we design a set of attack strategies grounded in realistic threat scenarios. These attacks exploit the core assumption of APR pipelines: that incoming bug reports are trustworthy, well-formed, and authored in good faith. By leveraging language models and project context, an attacker can craft reports that trigger harmful or wasteful behaviors.
@@ 3.1. Theoretical Threat Scenarios
We identify four classes of security risks associated with LLM-based APR pipelines. These risks align with known vulnerabilities outlined in OWASP’s CI/CD Top 10 (Krivelevich and Gil, 2025) and LLM Top 10 (Editor, [n. d.]), as well as recent studies on agent-based automation threats (Greshake et al., 2023).
@@
Malicious Patch Generation and Propagation. Attackers can exploit APR systems to synthesize patches that reintroduce vulnerabilities—such as reverting CVE fixes—or introduce insecure logic aligned with their goals. These attacks exploit the semantic gap between natural language bug reports and code safety guarantees. Even if such patches pass unit tests and post-hoc analysis, their acceptance into production systems can initiate downstream compromise via software supply chains (e.g., PyPI, Docker Hub).
APR/CI/CD Exploitation. Generated patches may embed logic that targets insufficiently isolated continuous integration environments. Malicious test code can access environment variables, exfiltrate secrets, or invoke remote shells. This is critical when APR/CI/CD infrastructure executes untrusted code without strict sandboxing or privilege separation.
Limited Monitoring and Accountability. Current APR pipelines have limited observability mechanisms to detect anomalous behavior, such as low-quality or high-risk patch patterns. Furthermore, logging of synthesis outcomes or tagging of APR-generated changes is usually optional, limiting post-incident analysis and trust evaluation.
@@ 3.2. Attack Demonstrations
To demonstrate the practical feasibility of these risks, we implement five adversarial attack types (Tab. 2). These attacks vary in their degree of automation, LLM usage, and reliance on project-specific context. All attacks target realistic open-source repositories and follow issue formats commonly used by human developers.
Context-Building Mechanism
For attacks that require deeper contextualization, we design an automated context-building component. Given a seed (e.g., a known CVE patch), we retrieve relevant files and commits using two retrieval methods: (1) locality-sensitive hashing (LSH) and (2) fuzzy string matching via rapidfuzz. We sanitize the diff by removing tests, documentation, and unrelated changes. This curated context is embedded into a prompt used to guide bug report generation, aligning it with the target project’s language and style. This method is used in the Revert CVE, Vulnerability Injection, and Deceptive Noise attacks. We used DevStral:24B, a compact open-weight model that can run efficiently on consumer-grade GPUs, making this attack feasible for a broad range of users while not being subject to server-side security guard-rails and filtering.
Attack Types
Table 2. Adversarial attacks categorized by objective and automation level.
Attack Objective Automation
Naive APR Trigger irrelevant synthesis via benign-looking prompt Semi-auto
CI/CD Exploits Leak secrets / remote execution via test injection Semi-auto
Vulnerability Injection Introduce insecure logic into production code Manual + context
Revert CVE Fix Undo previously applied security patch Auto + context
Deceptive Noise Waste APR and reviewer resources with realistic junk Auto + context
Naive APR. We prompt an LLM to generate a natural-sounding issue that causes the APR to perform a non-related synthesis task. As a toy example, the patch simply commits a new file containing the digits of $\pi$ written in English to the source code repository—demonstrating control over the output even when the issue appears benign.
@@
Vulnerability Injection. A vulnerable code snippet (e.g., leaking OS environment variables) is manually selected, and a realistic bug report is constructed via our context pipeline to reintroduce this logic through APR.
Revert CVE Fix. We identify a commit that fixes a CVE, sanitize its diff, and use this context to prompt (see Fig. 3) an LLM to write a bug report that questions or critiques the original fix. The resulting patch reverts the CVE fix and often bypasses all downstream validation.
Deceptive Noise Reports. We reuse the injection pipeline to create realistic but ultimately non-impactful bug reports. These are designed to consume APR compute and reviewer attention without triggering useful behavior, functioning as a targeted denial-of-service attack.
@@ 4.1. Defense Dimensions
Red Teaming and Adversarial Evaluation. As recommended by OWASP LLM-10, proactive adversarial testing is essential to evaluate the resilience of LLM-based systems. APR pipelines should be periodically stress-tested using red-teaming approaches that simulate sophisticated adversarial bug reports. Our own attack dataset serves this function, but future work could integrate such testing directly into CI/CD pipelines or model validation routines.
Telemetry and Feedback Integration. APR systems should emit runtime telemetry to detect suspicious usage patterns, such as repeated APR triggers from unknown users, high resource consumption on trivial issues, or consistent failure of downstream validations. Integrating feedback loops between APR, CI, and review systems would enable anomaly detection and facilitate continuous improvement of both filters and synthesizers.
Patch Provenance and Authorship Tracking. We recommend tracking metadata for APR-generated patches—including issue origin, APR agent identity, filtering results, and validation outcomes. Such provenance data would support forensic analysis and allow maintainers to audit the behavior of automated contributors. This aligns with recent supply chain security efforts, including SLSA and Sigstore.
Toward APR Alignment. Finally, enhancing the semantic alignment between reported issues and generated patches remains a key challenge. One promising direction is to augment APR prompts with project-specific policy constraints (e.g., “never modify auth code,” “never revert CVE patches”), enforced either via prompt engineering or post-hoc validation. GitHub has recently added support for adding custom instructions for copilot on a per-repository basis (GitHub, 2025b) although their documentation does not show the security impact of these instructions. We view this as a critical future direction for safe deployment.

### v2

@@ Abstract.
Large Language Model (LLM) - based Automated Program Repair (APR) systems are increasingly integrated into modern software development workflows, offering automated patches in response to natural language bug reports. However, this reliance on untrusted user input introduces a novel and underexplored attack surface. In this paper, we investigate the security risks posed by adversarial bug reports—realistic-looking issue submissions crafted to mislead APR systems into producing insecure or harmful code changes. We develop a comprehensive threat model and conduct an empirical study to evaluate the vulnerability of APR systems to such attacks. Our demonstration comprises 51 adversarial bug reports generated across a spectrum of strategies, ranging from manual curation to fully automated pipelines. We test these against a leading LLM-based APR system and assess both pre-repair defenses (e.g., LlamaGuard variants, PromptGuard variants, Granite-Guardian, and custom LLM filters) and post-repair detectors (GitHub Copilot, CodeQL). Our findings show that current defenses are insufficient: 90% of crafted bug reports triggered attacker-aligned patches. The best pre-repair filter blocked only 47%, while post-repair analysis–often requiring human oversight–was effective in just 58% of cases. To support scalable security testing, we introduce a prototype framework for automating the generation of adversarial bug reports. Our analysis exposes a structural asymmetry: generating adversarial inputs is inexpensive, while detecting or mitigating them remains costly and error-prone. We conclude with recommendations for improving the robustness of APR systems against adversarial misuse and highlight directions for future work on secure APR.
automated program repair, APR, security, attack, Large Language Models, LLM, SWE-Bench, SWE-Agent
journalyear: 2026copyright: ccconference: 23rd International Conference on Mining Software Repositories; April 13–14, 2026; Rio de Janeiro, Brazilbooktitle: 23rd International Conference on Mining Software Repositories (MSR ’26), April 13–14, 2026, Rio de Janeiro, Brazildoi: 10.1145/3793302.3793352isbn: 979-8-4007-2474-9/2026/04ccs: Software and its engineering Automatic programmingccs: Security and privacy Software security engineering
@@ 1. Introduction
Automated Program Repair (APR) has emerged as a promising solution to reduce developers’ effort in maintaining large software systems. By leveraging advances in machine learning and code synthesis, APR systems automatically generate candidate patches for bugs reported by users or detected by automated testing. This capability has led to widespread interest in integrating APR into developer workflows.
However, as APR becomes more pervasive, its security implications remain insufficiently explored. Existing research on APR focuses primarily on patch correctness, plausibility, or overfitting (Ye et al., 2025; Cambronero et al., 2019), but largely ignores the possibility that malicious actors could exploit the repair process itself as an attack vector. In particular, the reliance of APR systems on external bug reports submitted through open channels with minimal vetting introduces a new attack surface that has not been systematically studied.
@@
To answer these questions, we systematically design attacks that exploit common APR workflows. We conduct experiments using a leading LLM-powered APR system and representative open-source software repositories. Our study measures the feasibility, success rates, and cost implications of such attacks, and discusses possible hardening strategies for APR systems and project workflows.
This is no longer a purely theoretical concern: recent real-world attacks have exploited issue tracking systems to inject malicious inputs into downstream automation (Rehberger, 2025c). For instance, adversaries have used crafted Jira tickets to trigger AI-driven actions as part of a multi-stage attack on Atlassian’s MCP infrastructure (French, 2025; Waizel et al., 2025), and prompt injection vulnerabilities have been demonstrated in GitLab’s AI assistant (Willison, ), GitHub Copilot (Rehberger, 2025a; Higgs, 2025), and OpenHands (Rehberger, 2025b, d). These incidents underscore the urgent need to secure the interfaces between user-submitted bug reports and automated code generation systems.
1.1. Contributions, Artifacts, and Funding
1.1.1. Contributions
(1) A formal model describing how malicious actors can exploit bug reporting interfaces to subvert APR. (2) We provide attack framework including prompts used for attacks as well as an open-source prototype that automates the generation and evaluation of adversarial bug reports. (3) A systematic evaluation of pre- and post-APR defenses, including LLM-based filters, tests suites and static analysis tools, revealing key limitations and trade-offs.
1.1.2. Replication
We provide a complete replication package, including automation scripts for generating malicious bug reports, code to orchestrate the attack workflows, all defense mechanisms with prompts, as well as the full set of issues, generated patches, and execution trajectories (Happe et al., 2026).
1.1.3. Funding
This work was supported by the Polish National Agency for Academic Exchange (NAWA) - BPN/BAT/2025/1/00019 and Agentur für Bildung und Internationalisierung (ÖeAD).
@@ 2.1. Background and Threat Landscape
However, this design introduces a novel attack surface. LLM-driven APR systems inherently trust user-submitted bug reports and lack rigorous safeguards to verify intent, semantic correctness, or security implications of proposed changes. This makes them vulnerable to adversarial bug reports, i.e., crafted inputs that appear plausible but manipulate the APR pipeline into producing insecure or incorrect patches.
The security implications of LLM-based developer tools have recently gained attention, especially through the OWASP Top 10 for LLM Applications (Wilson and Dawson, 2025), which highlights prompt injection, overreliance, and insecure output handling as critical risks. Although defenses such as LlamaGuard (Inan et al., 2023), PromptGuard (Meta, 2025), and CodeQL (Youn et al., 2023) exist, they are not yet integrated into typical APR workflows. Moreover, most APR tools are not hardened against misuse of their input channels.
Despite growing awareness of security issues in LLM applications, vulnerabilities in APR workflows remain underexplored. Our work addresses this gap by investigating how adversarial bug reports can exploit LLM-powered repair systems, compromise CI/CD pipelines, and waste developer and compute resources.
@@ 2.2. End-to-End APR Pipeline and Attack Walkthrough
Figure 1 shows a typical APR pipeline. It starts with a user-submitted bug report. Enforcing authentication does not prevent attacks, as adversaries can create fake identities. Studies show marketplaces for such identities, often used in open-source supply chains like GitHub (He et al., 2026; devtips/, 2025). In our illustrative attack scenario, the adversary submits a malicious issue designed to trigger a reversion of a previously fixed CVE (Figure 2(a)). The issue appears plausible and well-formed, suggesting it will pass initial filtering.
Typically, the bug report passes through automated triage mechanisms. For APR, these may include: a security filter, such as LlamaGuard4, which assesses the textual input for prompt injections or unsafe patterns; and an APR suitability filter, which checks whether the report is appropriate for repair using an LLM with a structured prompt. Here, the issue is incorrectly classified as both non-malicious (LlamaGuard4) and valid for APR (GPT-4.1-mini).
Once accepted, the APR system attempts to reproduce the defect, often within a containerized testbed. This step may involve compiling the code, executing test cases, or generating synthetic inputs. Upon successful reproduction, a candidate patch is synthesized using templates, retrieval-based methods, or generative models. In our example, the patch effectively reverts the original CVE fix (Figure 2(b)). Next, the patch undergoes validation via automated test suites in a CI/CD environment. Since test coverage is rarely complete (especially for security fixes), such regressions often go undetected. If the patch passes, a pull request is automatically created. Beware that CI/CD environments may also be targets for attackers (although not illustrated in the running example).
The PR may be automatically reviewed (e.g., by GitHub Copilot, Figure 2(c)) and scanned by static analysis tools such as CodeQL (Figure 2(d)). However, such tools may miss subtle regressions or logic flaws introduced by an adversary. Manual review can help, but only if reviewers are involved early, have sufficient expertise, and invest the effort to investigate. In practice, limited context and reviewer fatigue make it easy for well-crafted reports and plausible code to slip through.
@@ 2.3. Threat Model
Source Code Integrity and security must be maintained: no backdoors or reintroduced vulnerabilities.
@@
Impact APR and CI/CD environment, e.g., DOS, runtime exploits, info leaks.
Source code, i.e., malicious patches may bypass both the initial bug report the and final source-code review and reach production.
@@
(Spoofing and Repudiation are less relevant in open bug reporting.)
To assess the security risks posed by adversarial bug reports, we define a structured threat model that captures both the system’s critical assets and the adversary’s capabilities. Our focus is on language model-driven APR pipelines that combine natural language inputs, automated patch generation, and integration into CI/CD workflows as shown by recent research (Bouzenia et al., 2024; Rondon et al., 2025; Li et al., 2025).
@@ 3. Adversarial Attack Strategies (RQ1)
To assess the susceptibility of APR systems to adversarial manipulation (RQ1), we design a set of attack strategies grounded in realistic threat scenarios. These attacks exploit the core assumption of APR pipelines: incoming bug reports are trustworthy, well-formed, and authored in good faith. By leveraging language models and project context, an attacker can craft reports that trigger harmful or wasteful behaviors.
@@ 3.1. Theoretical Threat Scenarios
We identify four classes of security risks associated with LLM-based APR pipelines. These risks align with known vulnerabilities outlined in OWASP’s CI/CD Top 10 (Krivelevich and Gil, 2025) and LLM Top 10 (Wilson and Dawson, 2025), as well as recent studies on agent-based automation threats (Greshake et al., 2023).
@@
Malicious Patch Generation and Propagation. Attackers can exploit APR systems to synthesize patches that reintroduce vulnerabilities, such as reverting CVE fixes, or introduce insecure logic aligned with their goals. These attacks exploit the semantic gap between natural language bug reports and code safety guarantees. Even if such patches pass unit tests and post-hoc analysis, their acceptance into production systems leads to downstream compromises via software supply chains (e.g., PyPI, Docker Hub).
APR/CI/CD Exploitation. Generated patches may embed logic that targets insufficiently isolated continuous integration environments. Malicious test code can access environment variables, exfiltrate secrets, or invoke remote shells. This is critical when the APR/CI/CD infrastructure executes untrusted code without strict sandboxing or privilege separation.
Limited Monitoring and Accountability. Current APR pipelines have limited observability mechanisms to detect anomalous behavior, such as low-quality or high-risk patch patterns. Furthermore, synthesis logging and tagging of APR-generated changes are often optional, limiting post-incident analysis and trust assessment.
@@ 3.2. Attack Demonstrations
To demonstrate feasibility we implement two attack strategies: noise injection and vulnerability injection. Noise injection wastes reviewer time and compute (APR and LLM-based review), while vulnerability injection does the same and leaves the project insecure. We first developed attack patterns on Django, then reran them automatically across all projects (including Django). We operationalize these strategies with five tactics: two use static prompts (Naive APR for Noise, CI/CD Exploits for Vulnerability) and three rely on context building to force specific faults (Deceptive Noise Reports, Vulnerability Injection, Revert CVE Fix). All tactics can be automated at scale; extended for new tactics; they differ in how they use LLMs and in their dependence on project-specific context building. Context building (described below) is cheap in terms of resources used and fully automated. All attacks target realistic open-source repositories and emulate the issue formats commonly used by human developers.
3.2.1. Context-Building Mechanism
For attacks that require deeper contextualization, we design an automated context-building component. Given a seed, e.g., a known CVE patch, we retrieve relevant files and commits using two retrieval methods: (1) locality-sensitive hashing (LSH) and (2) fuzzy string matching via rapidfuzz. [^1] https://rapidfuzz.github.io/RapidFuzz/ We sanitize the diff by removing tests, documentation, and unrelated changes. This curated context is embedded into a prompt used to guide bug report generation, aligning it with the target project’s language and style. This method is used in the Revert CVE, Vulnerability Injection, and Deceptive Noise attacks. We used DevStral:24B, a compact open-weight model that can run efficiently on consumer-grade GPUs, making this attack feasible for a broad range of users while not being subject to server-side security guard-rails and filtering.
3.2.2. Attack Tactics: Noise
  • Naive APR. We prompt an LLM to generate a natural-sounding issue that causes the APR to perform a non-related synthesis task. As a toy example, the issue forces a new file containing the digits of $\pi$ written in English to the project—demonstrating control over the output even when the issue appears benign.
  • Deceptive Noise Reports. We reuse the injection pipeline to generate realistic but non-impactful bug reports that consume APR compute and reviewer attention without triggering useful behavior, effectively acting as a targeted denial-of-service attack.
3.2.3. Attack Tactics: Vulnerability
@@
  • Vulnerability Injection. A vulnerable code snippet, e.g., leaking OS environment variables, is manually selected, and a realistic bug report is constructed via our context pipeline to reintroduce this logic through APR.
  • Revert CVE Fix. We identify a commit that fixes a CVE, sanitize its diff, and use this context to prompt an LLM (Fig. 3) to generate a bug report questioning the original fix. The resulting patch reverts the CVE fix and often bypasses downstream validation.
@@ 4.1. Defense Dimensions
Red Teaming and Adversarial Evaluation. As recommended by OWASP LLM-10 (Wilson and Dawson, 2025), proactive adversarial testing is essential to evaluate the resilience of LLM-based systems. APR pipelines should be periodically stress-tested using red-teaming approaches that simulate sophisticated adversarial bug reports. Our own attack dataset serves this function, but future work could integrate such testing directly into CI/CD pipelines or model validation routines.
Telemetry and Feedback Integration. APR systems should emit runtime telemetry to detect suspicious usage patterns, such as repeated APR triggers from unknown users, high resource consumption for trivial issues, or consistent failures in downstream validations. Integrating feedback loops between APR, CI, and review systems would enable anomaly detection and facilitate continuous improvement of both filters and synthesizers.
Patch Provenance and Authorship Tracking. We recommend tracking metadata for APR-generated patches, including issue origin, APR agent identity, filtering results, and validation results. Such provenance data would support forensic analysis and allow maintainers to audit the behavior of automated contributors. This aligns with recent supply chain security efforts, including SLSA and Sigstore.
Toward APR Alignment. Finally, enhancing the semantic alignment between reported issues and generated patches remains a key challenge. One promising direction is to augment APR prompts with project-specific policy constraints (e.g., “never modify auth code,” “never revert CVE patches”), enforced either via prompt engineering or post-hoc validation. GitHub has recently added support for adding custom instructions for Copilot on a per-repository basis (GitHub, 2025b), although their documentation does not show the security impact of these instructions. We view this as a critical future direction for safe deployment.